What would be the coefficient of determination if the total sum of squares (SST) is 23.29 and the sum of squares due to regression (SSR) is 10.03
Answer #1

From the given information, we have \(S S R=10.03, S S T=23.29\) The coefficient of determination is equal to the regression sum of squares (that is, explained variation) divided by the total sum of squares (that is, total variation). \(r^{2}=\frac{\text { Regression sum of squares }}{\text { Total sum of squares }}\) \(=\frac{S S R}{S S T}\)

\(=\frac{10.03}{23.29}\)

\(=0.430657\)

\(r^{2}=43.06 \%\)

Interpretation:

The value of the coefficient of determination \(r^{2}=43.06 \%\) percentage of the variation in the dependent variable \((Y)\) can be explained by the independent variable \((X)\)